OpenStudy (anonymous):

the length of a rectangular storage shed must be 4 meters more than the width, and the perimeter must be at least 120 meters. What is the range of values for the width?

OpenStudy (anonymous):

l=w+4 P=2w+2l P=2w+2(w+4) P=2w+2w+8 P=4w+8 P>120, therefore 4w+8>120 4w>112 w>28 The width must be greater than 28 meters.

OpenStudy (anonymous):

thank you. i have one more that is a long one.

OpenStudy (anonymous):

To clarify, the width must be greater than *or equal to* 28 meters, since the perimeter must be greater than or equal to 120 meters.